Suppose that quiz scores in a beginning statistics class have a mean of 7.2 with a standard deviation of 0.4. Using Chebyshev's Theorem, state the range in which at least 75% of the data will reside. Please do not round your answers.
Answer #1

Answer:

Given,

Mean = 7.2 , Standard deviation = 0.4

consider,

1 - 1/k^2 = 0.75

1/k^2 = 0.25

k^2 = 1/0.25

k^2 = 4

k = 2

Now consider,

Range = mean +/- 2*sd

= 7.2 +/- 2*0.4

= 7.2 +/- 0.8

= (6.4 , 8)